# Categorical Reparameterization with Denoising Diffusion models

### v1

@@ Abstract
Gradient-based optimization with categorical variables typically relies on score-function estimators, which are unbiased but noisy, or on continuous relaxations that replace the discrete distribution with a smooth surrogate admitting a pathwise (reparameterized) gradient, at the cost of optimizing a biased, temperature-dependent objective. In this paper, we extend this family of relaxations by introducing a diffusion-based soft reparameterization for categorical distributions. For these distributions, the denoiser under a Gaussian noising process admits a closed form and can be computed efficiently, yielding a training-free diffusion sampler through which we can backpropagate. Our experiments show that the proposed reparameterization trick yields competitive or improved optimization performance on various benchmarks.
@@ 1 Introduction
Many learning problems involve discrete choices– actions in reinforcement learning, categorical latent variables in variational inference, token-level decisions in sequence modeling, or combinatorial assignments in structured prediction and discrete optimization. A common primitive is the minimization of an objective of the form $\mathbb{E}_{\pi_{\!\theta}}[f(X)]$, where $\pi_{\!\theta}$ is a categorical distribution, often with a mean-field structure, over $L$ discrete variables, each taking values in a vocabulary of size $K$. The function $f$ represents a downstream loss or constraint penalty evaluated on discrete samples, typically through one-hot encodings. Computing $\nabla_{\!{\theta}}\mathbb{E}_{\pi_{\!\theta}}[f(X)]$ exactly is generally intractable: in the absence of exploitable structure in $f$, it requires summing over $K^{L}$ configurations. The challenge, therefore, is to construct gradient estimators that are both computationally feasible and have a low mean squared error.
Existing estimators exhibit a standard bias–variance trade-off. Score-function estimators, such as REINFORCE (williams1992simple; greensmith2004variance), are unbiased but often suffer from high variance, which motivates the use of variance-reduction techniques most often using learned control variates (tucker2017rebar; grathwohl2018relax); they often yield useful gradients in practice but are biased with respect to the true discrete objective, with recent refinements such as ReinMax improving the approximation (liu2023reinmax). Continuous relaxations based on approximate reparameterizations, most notably the Gumbel-Softmax / Concrete construction (maddison2016concrete; jang2017categorical), replace $\pi_{\!\theta}$ by a smooth family on the simplex controlled by a temperature parameter. While this enables pathwise differentiation, taking the temperature small to reduce bias drives the sampler towards an argmax map and typically leads to ill-conditioned or vanishing gradients, whereas larger temperatures optimize a substantially different objective.
In this work, we revisit continuous relaxations through the lens of denoising diffusion models (sohl2015deep; song2019generative; ho2020denoising). Diffusion models generate data by transforming a Gaussian sample into a sample from the target data distribution through iterative denoising dynamics, which are explicitly constructed as the reverse of a chosen forward noising process. In practice, implementing the sampler requires only access to a denoiser; that is, a function that, given a noisy input and its noise level or time index, returns the expected clean signal.
@@
In this paper, we exploit the key observation that for a categorical distribution supported on simplex vertices, the denoiser at each noise level can be computed in closed form. This enables us to construct a training-free, diffusion-based, differentiable, and approximate sampling map from Gaussian noise to the categorical distribution $\pi_{\!\theta}$. We then analyze the small-noise regime, which serves as the temperature parameter, and characterize the emergence of nearly constant transport regions and sharp decision boundaries, explaining when and why gradients become uninformative as the relaxation approaches the discrete target.We derive practical gradient estimators, including hard variants, that recover the hard Straight-Through (bengio2012estimating) and ReinMax (liu2023reinmax) as special cases when using a single diffusion step. We also propose a parameter-dependent initialization that improves performance while keeping the diffusion overhead small. Empirically, preliminary experiments show that our approach yields competitive or improved optimization performance on various benchmarks.
@@
We denote the $K$-simplex by $\Delta^{K-1}$. For a matrix $x\in{\mathbb{R}}^{L\times K}$, we write $x^{i}\in{\mathbb{R}}^{K}$ for its $i$-th row and $x^{ij}$ for the $(i,j)$-th element. The softmax operator on a matrix $x\in{\mathbb{R}}^{L\times K}$ is defined row-wise by $\text{softmax}(x)\in{\mathbb{R}}^{L\times K}$ with entries $\text{softmax}(x)^{ik}=\exp(x^{ik})/\sum_{j=1}^{K}\exp(x^{ij})$ for $(i,k)\in[L]\times[K]$. For a map $f:{\mathbb{R}}^{d}\to{\mathbb{R}}^{m}$, we write $\operatorname{J}_{x}\!f\in{\mathbb{R}}^{m\times d}$ for its Jacobian matrix. To write Jacobians for maps $f:{\mathbb{R}}^{L^{\prime}\times K^{\prime}}\to{\mathbb{R}}^{L\times K}$ conveniently, we implicitly identify matrices with their vectorized forms, obtained by stacking all rows into a single column vector. Gradients and Jacobians are taken with respect to these vectorized representations, and we do not distinguish notationally between a matrix and its vectorization.
@@ 2 Background
where $f:\mathsf{X}\times\Theta\to{\mathbb{R}}$, $\Theta\subseteq{\mathbb{R}}^{m}$, and $\{\pi_{\theta}:\theta\in\Theta\}$ is the parameterized family of probability mass functions (p.m.f.) over $\mathsf{X}$. Without loss of generality, we assume that $\mathsf{X}=\mathsf{V}^{L}$ for some $L\in\mathbb{N}$, where $\mathsf{V}\coloneqq\{e_{k}\}_{k=1}^{K}$ denotes the set of $K$ one-hot encodings, and $e_{k}$ is the one-hot vector with $1$ at position $k$.We also assume that the distribution $\pi_{\!\theta}$ factorizes according to this categorical structure: for any $\theta\in\Theta$ and $x=(x^{1},\ldots,x^{L})\in\mathsf{X}$,
$\pi_{\!\theta}(x)=\prod_{i=1}^{L}\pi^{i}_{\!\theta}(x^{i}),\;\pi^{i}_{\!\theta}(x^{i})\coloneqq\frac{\exp(\langle x^{i},\varphi_{\theta}^{i}\rangle)}{\sum_{j=1}^{K}\exp(\langle x^{j},\varphi_{\theta}^{i}\rangle)}\;,\!\!$ (2)
where $\theta\mapsto\varphi_{\theta}\in{\mathbb{R}}^{L\times K}$ is such that $\varphi^{i}_{\theta}$ are the logits of the $i$-th categorical component. The factorization (2) is standard and is used in reinforcement learning to model policies (wu2018variance; berner2019dota; vinyals2019alphastar), in training Boltzmann machines (hinton2012practical), in VQ-VAEs (van2017neural), and more recently for modeling transitions in discrete diffusion models.(hoogeboom2021argmax; austin2021structured; campbell2022continuous; lou2023discrete; shi2024simplified; sahoo2024simple).
Under mild regularity assumptions on $f$ and $(\varphi^{i}_{\theta})_{i=1}^{L}$, the gradient of (1) is given by
$\nabla_{\!{\theta}}F(\theta)=\mathbb{E}_{\pi_{\!\theta}}[\nabla_{\!{\theta}}f_{\theta}(X)]+\sum_{x}f_{\theta}(x)\nabla_{\!{\theta}}\pi_{\!\theta}(x)$ (3)
and is intractable as the sum ranges over $K^{L}$ states. Nevertheless, in the case where $f$ is separable across the dimensions, i.e., $f_{\theta}(x)=\sum_{i}^{L}f^{i}_{\theta}(x^{i})$, for a family of functions $\{f^{i}:\mathsf{V}\times\Theta\to{\mathbb{R}}\}_{i=1}^{L}$, as is the case for many information-theoretic divergences, the expectation separates into $L$ independent terms, reducing the computation to $\mathcal{O}(LK)$. Furthermore, while the first term can be approximated via Monte Carlo, the second term can also be estimated similarly using the REINFORCE identity (williams1992simple),
$\sum_{x}f_{\theta}(x)\nabla_{\!{\theta}}\pi_{\!\theta}(x)=\mathbb{E}_{\pi_{\!\theta}}[f_{\theta}(X)\nabla_{\!{\theta}}\log\pi_{\!\theta}(X)]\;.$
It is well known, however, that the vanilla Monte Carlo estimator of the r.h.s. suffers from high variance (sutton_barto_2018). In practice, it is used together with baselines or other control-variate techniques to reduce variance (greensmith2004variance; mnih2014neural; mnih2016variational; tucker2017rebar; titsias2022double; grathwohl2018relax). Other estimation methods have been proposed, such as the Straight-Through estimator (bengio2012estimating) or Gumbel-Softmax reparameterization (maddison2016concrete; jang2017categorical), which we now review. For ease of presentation and without loss generality, we assume in the remainder of the paper that $f$ does not depend on $\theta$, i.e., $f_{\theta}(x)=f(x)$.
@@
Popular estimators either replace the objective $F$ by a differentiable surrogate and use its gradient, or directly construct a surrogate for $\nabla_{\!{\theta}}F$ itself. From now on we assume that $f$ is differentiable w.r.t. $x$. One such estimator is the Straight-Through (ST) approach, which replaces the discrete objective by the surrogate obtained by swapping $f$ and the expectation in (1), and differentiates the map $\theta\mapsto f\!\bigl(\mathbb{E}_{\pi_{\!\theta}}[X]\bigr)$. Noting that the Jacobian of $\varphi_{\theta}\mapsto\mathbb{E}_{\pi_{\!\theta}}[X]$ is $\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)\in{\mathbb{R}}^{LK\times LK}$, the gradient of this surrogate w.r.t. the logits is $\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)\nabla_{\!{x}}f(\mathbb{E}_{\pi_{\!\theta}}[X]).$ A popular practical instance of ST replaces the expectation inside $\nabla_{\!{x}}f$ with a single Monte Carlo sample $X\sim\pi_{\!\theta}$, often referred to as hard ST;
@@
This gradient estimator was first considered by hinton2012nnml_lec9 in the context of training with hard thresholds, where the backward pass treats the threshold operation as the identity. It was later formalized by bengio2012estimating for quantization-aware training of deep networks. The resulting gradient estimator is often effective in practice but is, by construction, biased with respect to the true discrete objective. When $f$ is linear, hard ST yields an unbiased gradient of $F$. As observed by liu2023reinmax, the hard ST surrogate can be interpreted as an unbiased estimator of a first-order approximation of $\nabla_{\!{\theta}}F(\theta)$ defined in Appendix B.
The ReinMax estimator (liu2023reinmax) improves upon hard ST by using an unbiased estimator of a second-order approximation of $\nabla_{\!{\theta}}F(\theta)$, obtained via the trapezoidal rule. In Appendix B, we prove that the resulting estimator has the following simple form, which closely resembles hard ST:
$\widehat{\nabla}^{\text{RM}}_{\!\theta}F(X;\theta)\\ \coloneqq\frac{1}{2}\operatorname{J}_{\theta}\!\varphi_{\theta}\!{}^{\top}\!\big\{\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)+\widehat{C}_{\theta}(X)\big\}\nabla_{\!{x}}f(X)$ (5)
where $X\sim\pi_{\!\theta}$ and $\widehat{C}_{\theta}(X)\coloneqq(X-\mathbb{E}_{\pi_{\!\theta}}[X])(X-\mathbb{E}_{\pi_{\!\theta}}[X])^{\top}$ is an unbiased estimator of $\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)$. While hard ST is exact for linear $f$, ReinMax is exact for quadratic $f$. A more detailed discussion of ReinMax is provided in Appendix B.
@@
A second family of gradient estimators relies on approximate reparameterization techniques for discrete random variables. The reparameterization trick was originally introduced for continuous distributions, where a sample can be expressed as a deterministic transformation of an auxiliary latent variable (kingma2013auto). Specifically, we temporarily assume that $\pi_{\!\theta}$ is a distribution that admits a reparameterization, that is,
$\pi_{\!\theta}\coloneqq\text{Law}\big(T_{\theta}(Z)\big),$ (6)
where $Z\sim p$ with $p$ a distribution that does not depend on $\theta$, and $T_{\theta}$ is a measurable transformation mapping $Z$ to a sample distributed according to $\pi_{\!\theta}$. Assume also that for $p$-almost every $z$ the map $\theta\mapsto T_{\theta}(z)$ is differentiable for any $\theta\in\Theta$ and that $z\mapsto\nabla_{\!{\theta}}f(T_{\theta}(z))$ satisfies standard domination conditions for all $\theta\in\Theta$, so that differentiation under the expectation is justified by the Lebesgue dominated convergence theorem. Then
@@
which yields a low-variance Monte Carlo estimator of the objective gradient (schulman2015gradient). Following previous works with refer to such estimators as pathwise or reparameterized gradient.
In the discrete case however, such an exact reparameterization is not available. Indeed, any representation of $\pi_{\!\theta}$ as the pushforward of a simple continuous base distribution typically yields a map $\theta\mapsto T_{\theta}(z)$ that is piecewise constant with jump discontinuities. In particular, $\operatorname{J}_{\theta}\!T_{\theta}(z)=0$ for almost every $(z,\theta)$ and so $\mathbb{E}[\operatorname{J}_{\theta}\!T_{\theta}(Z)]=0$ while $\nabla_{\!{\theta}}F(\theta)\neq 0$, illustrating that (7) does not hold in this discrete setting because the differentiability at every $\theta\in\Theta$ breaks and the domination requirement for swapping limit and integral is violated. As a simple example, consider $\pi_{\!\theta}=\text{Bernoulli}(\theta)$ and $f(x)=x$. A possible choice of transform is $T_{\theta}(Z)=\mathds{1}_{\left(-\infty,\theta\right]}(Z)$ with $Z\sim\text{Uniform}([0,1])$ which yields to zero gradient almost surely and therefore (7) does not hold. To circumvent this problem, one typically uses continuous relaxations of $\pi_{\!\theta}$ for which (7) is valid and that trade bias for lower-variance gradients
The Gumbel–Softmax (or Concrete) distribution (maddison2016concrete; jang2017categorical) is a canonical example of such a relaxation: it replaces the categorical distribution $\pi_{\!\theta}$ (on the edges of the simplex) with a temperature-indexed family of continuous distributions $(\pi^{\theta}_{\tau})_{\tau>0}$ on the simplex, each of which admits a pathwise gradient estimator satisfying (7). Specifically, $\pi^{\theta}_{\tau}\;\coloneqq\;\text{Law}\big(T^{\theta}_{\tau}(G)\big)$ is used as a relaxed surrogate for $\pi_{\!\theta}$, where for all $\theta\in\Theta$,
$T^{\theta}_{\tau}(G)\;\coloneqq\;\text{softmax}\big((\varphi_{\theta}+G)/\tau\big)\;,\quad\tau>0$
and $G\in{\mathbb{R}}^{L\times K}$ is a random matrix with i.i.d. Gumbel entries $G^{ij}\sim\mathrm{Gumbel}(0,1)$. As $\tau\to 0$, $\{\pi^{\theta}_{\tau}:\,\tau>0\}$ converges in distribution to the distribution of the random matrix $T_{\theta}(G)\coloneqq\operatorname*{argmax}_{x\in\mathsf{X}}(\varphi_{\theta}+G)x^{\top}$ which is precisely $\pi_{\!\theta}$; see gumbel1954statistical. This is known as the Gumbel-max trick. It is easy to verify that for the surrogate objective $F_{\tau}(\theta)=\mathbb{E}_{\pi^{\theta}_{\tau}}[f(X)]$, which converges to $F$ as $\tau\to 0$, (7) holds under appropriate assumptions on $f$, thus allowing an approximate reparameterization trick at the expense of a certain bias controlled by the parameter $\tau$.

### v2

@@ Abstract
Learning models with categorical variables requires optimizing expectations over discrete distributions, a setting in which stochastic gradient-based optimization is challenging due to the non-differentiability of categorical sampling. A common workaround is to replace the discrete distribution with a continuous relaxation, yielding a smooth surrogate that admits reparameterized gradient estimates via the reparameterization trick. Building on this idea, we introduce ReDGE, a novel and efficient diffusion-based soft reparameterization method for categorical distributions. Our approach defines a flexible class of gradient estimators that includes the Straight-Through estimator as a special case. Experiments spanning latent variable models and inference-time reward guidance in discrete diffusion models demonstrate that ReDGE consistently matches or outperforms existing gradient-based methods. The code is available at https://github.com/samsongourevitch/redge.
@@ 1 Introduction
Many learning problems involve discrete choices, such as actions in reinforcement learning, categorical latent variables in variational inference, token-level decisions in sequence modeling, or combinatorial assignments in structured prediction and discrete optimization. A common primitive in these settings is the minimization of an objective of the form $\mathbb{E}_{\pi_{\!\theta}}[f(X)]$, where $\pi_{\!\theta}$ is a categorical distribution corresponding to the law of $L$ independent discrete random variables each taking values in a vocabulary of size $K$. The function $f$ represents a downstream loss or constraint penalty evaluated on discrete samples, typically through one-hot encodings. Computing $\nabla_{\!{\theta}}\mathbb{E}_{\pi_{\!\theta}}[f(X)]$ exactly is generally intractable: in the absence of exploitable structure in $f$, it requires summing over $K^{L}$ configurations. The challenge, therefore, is to construct gradient estimators that are both computationally feasible and have a low mean squared error.
Existing estimators exhibit a standard bias–variance trade-off. Score-function estimators, such as Reinforce (Williams, 1992; Greensmith et al., 2004), are unbiased but often suffer from high variance, which motivates the use of variance-reduction techniques most often using learned control variates (Tucker et al., 2017; Grathwohl et al., 2018); they often yield useful gradients in practice but are biased with respect to the true discrete objective, with recent refinements such as ReinMax improving the approximation (Liu et al., 2023a). Continuous relaxations based on approximate reparameterizations, most notably the Gumbel-Softmax / Concrete construction (Maddison et al., 2017; Jang et al., 2017), replace $\pi_{\!\theta}$ by a smooth family on the simplex controlled by a temperature parameter. While this enables pathwise differentiation, taking the temperature small to reduce bias drives the sampler towards an argmax map and typically leads to ill-conditioned or vanishing gradients, while higher temperatures provide well-behaved gradients but correspond to optimizing a more relaxed objective.
In this work, we revisit continuous relaxations through the lens of denoising diffusion models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020). Diffusion models generate data by transforming a Gaussian sample into a sample from the target data distribution through iterative denoising dynamics explicitly constructed as the reverse of a chosen forward noising process. In practice, implementing the sampler requires only access to a denoiser; that is, a function that, given a noisy input and its noise level or time index, returns the expected clean signal.
@@
We exploit the key observation that for a categorical distribution supported on simplex vertices, the denoiser at each noise level can be computed in closed form. This enables us to construct a training-free, diffusion-based, differentiable, and approximate sampling map from Gaussian noise to the categorical distribution $\pi_{\!\theta}$. We then analyze the small-noise regime, which serves as the temperature parameter, and characterize the emergence of nearly constant transport regions and sharp decision boundaries, explaining when and why gradients become uninformative as the relaxation approaches the discrete target. We derive practical gradient estimators, including hard variants, that recover the hard Straight-Through (Bengio et al., 2013) and ReinMax (Liu et al., 2023a) as special cases when using a single diffusion step. We also propose a parameter-dependent initialization that improves performance while keeping the diffusion overhead small. Empirically, we find that our diffusion-based reparameterization yields strong results across a diverse set of benchmarks, including polynomial programming, variational inference, and inference-time reward guidance, typically matching or improving upon prior estimators.
@@
For any positive integer $n$, let $n]\coloneqq\{1,\ldots,n\}$. We denote the $K$-simplex by $\Delta^{K-1}=\{w\in{\mathbb{R}}^{K}_{+}\,:\,\sum_{k=1}^{K}w^{k}=1\}$. For a matrix $x\in{\mathbb{R}}^{L\times K}$, we write $x^{i}\in{\mathbb{R}}^{K}$ for its $i$-th row and $x^{ij}$ for the $(i,j)$-th element. Besides, we identify any $x\in{\mathbb{R}}^{L\times K}$ as a vector using the row-major order in which the matrix elements are ordered by row. The softmax operator on a matrix $x\in{\mathbb{R}}^{L\times K}$ is defined row-wise by $\text{softmax}(x)\in{\mathbb{R}}^{L\times K}$ with entries $\text{softmax}(x)^{ik}=\exp(x^{ik})/\sum_{j=1}^{K}\exp(x^{ij})$ for $(i,k)\in[L]\times[K]$. For a map $f:{\mathbb{R}}^{d}\to{\mathbb{R}}^{m}$, we write $\operatorname{J}_{x}\!f\in{\mathbb{R}}^{m\times d}$ for its Jacobian matrix. To write Jacobians for maps $f:{\mathbb{R}}^{L^{\prime}\times K^{\prime}}\to{\mathbb{R}}^{L\times K}$ conveniently, we implicitly identify matrices with their vectorized forms. Gradients and Jacobians are taken with respect to these vectorized representations, and we do not distinguish notationally between a matrix and its vectorization.
@@ 2 Background
where $f:\mathsf{X}\times\Theta\to{\mathbb{R}}$, $\Theta\subseteq{\mathbb{R}}^{m}$, and $\{\pi_{\theta}:\theta\in\Theta\}$ is a parameterized family of probability mass functions (p.m.f.) over $\mathsf{X}$. Without loss of generality, we assume that $\mathsf{X}=\mathsf{V}^{L}$ for some $L\in\mathbb{N}$, where $\mathsf{V}\coloneqq\{e_{k}\}_{k=1}^{K}$ denotes the set of $K$ one-hot encodings, and $e_{k}\in{\mathbb{R}}^{K}$ is the one-hot vector with $1$ at position $k$. We also assume that the distribution $\pi_{\!\theta}$ factorizes according to this categorical structure: for any $\theta\in\Theta$ and $x=(x^{1},\ldots,x^{L})\in\mathsf{X}$,
$\pi_{\!\theta}(x)=\prod_{i=1}^{L}\pi^{i}_{\!\theta}(x^{i}),\;\pi^{i}_{\!\theta}(x^{i})\coloneqq\frac{\exp(\langle x^{i}\;,\varphi_{\theta}^{i}\rangle)}{\sum_{j=1}^{K}\exp(\varphi_{\theta}^{ij})}\;,\!\!$ (2)
where $\theta\mapsto\varphi_{\theta}\in{\mathbb{R}}^{L\times K}$ is such that $\varphi^{i}_{\theta}$ are the logits of the $i$-th categorical component. The factorization (2) is standard and is used in reinforcement learning to model policies (Wu et al., 2018; Berner et al., 2019; Vinyals et al., 2019), in training Boltzmann machines (Hinton, 2012), in VQ-VAEs (Van Den Oord et al., 2017), and more recently for modelling transitions in discrete diffusion models.(Hoogeboom et al., 2021; Austin et al., 2021; Campbell et al., 2022; Lou et al., 2023; Shi et al., 2024; Sahoo et al., 2024).
Under mild regularity assumptions on $f$ and $\varphi_{\theta}$, the gradient of (1) is given by
$\nabla_{\!{\theta}}F(\theta)=\mathbb{E}_{\pi_{\!\theta}}[\nabla_{\!{\theta}}f_{\theta}(X)]+\sum_{x\in\mathsf{X}}f_{\theta}(x)\nabla_{\!{\theta}}\pi_{\!\theta}(x)$ (3)
and is intractable as the sum ranges over $K^{L}$ states. Furthermore, while the first term can be approximated via Monte Carlo, the second term has to be estimated separately. One option is to use the reinforce estimator. (Williams, 1992) However, it is well known that the vanilla forma of this estimator suffers from high variance (Sutton and Barto, 2018) and has be to combined with baselines or other control-variate techniques to reduce variance (Greensmith et al., 2004; Mnih and Gregor, 2014; Mnih and Rezende, 2016; Tucker et al., 2017; Titsias and Shi, 2022; Grathwohl et al., 2018). Other estimation methods have been proposed, such as the Straight-Through estimator (Bengio et al., 2013) or Gumbel-Softmax reparameterization (Maddison et al., 2017; Jang et al., 2017), which we briefly review here.
For simplicity, we assume throughout that $f$ does not depend on $\theta$ (i.e., $f_{\theta}(x)=f(x)$) and is differentiable in $x$.
@@
Popular estimators either replace the objective $F$ by a differentiable surrogate and use its gradient, or directly construct a surrogate for $\nabla_{\!{\theta}}F$ itself. One such estimator is the Straight-Through (ST) approach, which replaces the discrete objective by the surrogate obtained by swapping $f$ and the expectation in (1), and differentiates the map $\theta\mapsto f(\mathbb{E}_{\pi_{\!\theta}}[X])$. Noting that $\operatorname{J}_{\theta}\!\mathbb{E}_{\pi_{\!\theta}}[X]=\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)\operatorname{J}_{\theta}\!\varphi_{\theta}$, the gradient of this surrogate is $\operatorname{J}_{\theta}\!\varphi_{\theta}^{\top}\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)\nabla_{\!{x}}f(\mathbb{E}_{\pi_{\!\theta}}[X]).$ A popular practical instance of ST replaces the expectation inside $\nabla_{\!{x}}f$ with a single Monte Carlo sample $X\sim\pi_{\!\theta}$, often referred to as hard ST:
@@
This gradient estimator was first considered by Hinton et al. (2012) in the context of training with hard thresholds, where the backward pass treats the threshold operation as the identity. It was later formalized by Bengio et al. (2013) for quantization-aware training of deep networks. The resulting gradient estimator is often effective in practice but is, by construction, biased with respect to the true discrete objective. When $f$ is linear, hard ST yields an unbiased gradient of $F$.
The ReinMax estimator (Liu et al., 2023a) refines hard ST by providing an exact unbiased estimator of $\nabla_{\!{\theta}}F(\theta)$ in case $f$ is quadratic, obtained via a trapezoidal (Heun-type) rule. In Appendix B, we show that it admits the following simple form, closely mirroring hard ST:
$\small\widehat{\nabla}^{\text{RM}}_{\!\theta}F(X;\theta)\coloneqq\frac{1}{2}\,\operatorname{J}_{\theta}\!\varphi_{\theta}^{\top}B_{\theta}(X)\nabla_{\!{x}}f(X)\;,$ (5)
where $X\sim\pi_{\!\theta}$. Here $B_{\theta}(X)=\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)+\widehat{C}_{\theta}(X)$, where $\widehat{C}_{\theta}(X)$ is block-diagonal with $L$ blocks of size $K\times K$; its $\ell$-th block is $\widehat{C}^{(\ell)}_{\theta}(X)\coloneqq(X^{\ell}-\mathbb{E}_{\pi_{\!\theta}}[X^{\ell}])(X^{\ell}-\mathbb{E}_{\pi_{\!\theta}}[X^{\ell}])^{\top}$, implying that $\mathbb{E}_{\pi_{\!\theta}}[\widehat{C}^{(\ell)}_{\theta}(X)]=\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X^{\ell})$. We provide further details and proofs in Appendix B.
@@
For continuous distributions, the reparameterization trick expresses a sample as a deterministic transform of auxiliary noise (Kingma and Welling, 2013b). Specifically, we temporarily assume that $\pi_{\!\theta}$ is a distribution that admits a reparameterization, that is, $\pi_{\!\theta}\coloneqq\text{Law}\big(T_{\theta}(Z)\big),$ where $Z$ follows a distribution $p$ that does not depend on $\theta$, typically uniform or Gaussian. Assume also that for $p$-almost every $z$ the map $\theta\mapsto T_{\theta}(z)$ is differentiable for any $\theta\in\Theta$ and that $z\mapsto\nabla_{\!{\theta}}f(T_{\theta}(z))$ satisfies standard domination conditions for all $\theta\in\Theta$, so that differentiation under the expectation is justified by the Lebesgue dominated convergence theorem. Then
@@
which yields a low-variance Monte Carlo estimator of the objective gradient (Schulman et al., 2015). In the discrete case however, such an exact reparameterization is not available. Any representation of $\pi_{\!\theta}$ as the pushforward of a simple continuous base distribution typically yields a map $\theta\mapsto T_{\theta}(z)$ that is piecewise constant with jump discontinuities. As a consequence, for any $\theta$, $\operatorname{J}_{\theta}\!T_{\theta}(z)=0$ for almost every $z$, and therefore $\mathbb{E}[\operatorname{J}_{\theta}\!T_{\theta}(Z)^{\top}\nabla_{\!{x}}f(T_{\theta}(Z))]=0$ while $\nabla_{\!{\theta}}F(\theta)\neq 0$. Thus (6) does not hold in the discrete setting as the differentiability at every $\theta\in\Theta$ fails, and the domination condition needed to justify differentiation under the integral sign is violated. To circumvent this issue, one typically resorts to continuous relaxations of $\pi_{\!\theta}$, i.e., distributions admitting a density with respect to the Lebesgue measure. For such relaxations, (6) is valid, at the cost of introducing bias in exchange for lower-variance gradient estimates.
The Gumbel–Softmax (or Concrete) distribution (Maddison et al., 2017; Jang et al., 2017) is a canonical example of such a relaxation: $\pi_{\!\theta}$ is replaced with a temperature-indexed family of continuous distributions $(\pi^{\theta}_{\!\tau})_{\tau>0}$ on the simplex that admit pathwise gradient estimator satisfying (6). Specifically, $\pi^{\theta}_{\!\tau}\;\coloneqq\;\text{Law}\big(T^{\theta}_{\tau}(G)\big)$ is used as a relaxed surrogate for $\pi_{\!\theta}$, where for all $\theta\in\Theta$,
$T^{\theta}_{\tau}(G)\;\coloneqq\;\text{softmax}\big((\varphi_{\theta}+G)/\tau\big)\;,\quad\tau>0\;,$
and $G\in{\mathbb{R}}^{L\times K}$ is a random matrix with i.i.d. Gumbel entries $G^{ij}\sim\mathrm{Gumbel}(0,1)$. As $\tau\to 0$, $\{\pi^{\theta}_{\!\tau}:\,\tau>0\}$ converges in distribution to $\pi_{\!\theta}$; see Gumbel (1954). This is known as the Gumbel-max trick (Maddison et al., 2017). It is easy to verify that for the surrogate objective $F_{\tau}(\theta)=\mathbb{E}_{\pi^{\theta}_{\!\tau}}[f(X)]$, which converges to $F$ as $\tau\to 0$, (6) holds under appropriate assumptions on $f$, thus allowing an approximate reparameterization trick at the expense of a certain bias controlled by the parameter $\tau$.
Remark 1.
To compute the true gradient $\nabla_{\!{\theta}}F(\theta)$, only the values of $f$ on $\mathsf{X}$ are relevant. In contrast, the estimators considered here differentiate a fixed continuous extension of $f$ to $(\Delta^{K-1})^{L}$ (or to ${\mathbb{R}}^{L\times K}$) which we assume is provided by the downstream model and denote again by $f$. Note however that distinct extensions may agree on $\mathsf{X}$ while inducing different gradients on the simplex. Throughout, we avoid this ambiguity by treating the choice of extension as part of the problem specification.